OpenStudy (anonymous):

find the exact value of cos(pi/12) using sum and difference formulas

OpenStudy (anonymous):

\[\frac{1}{12}=\frac{1}{3}-\frac{1}{4}\] and so \[\frac{\pi}{12}=\frac{\pi}{3}-\frac{\pi}{4}\] use the "difference formula" for \[\cos(A-B)\] with \(A=\frac{\pi}{3}\) and \(B=\frac{\pi}{4}\)